OpenStudy (anonymous):

30. What is the volume of a square pyramid with base edges of 32cm and a slant height of 34?

OpenStudy (calculusxy):

\[\frac{ 1 }{ 3 }(B \times h)\]

OpenStudy (calculusxy):

\[B \] represents the area of the base (square): \[length \times width\]

OpenStudy (calculusxy):

http://www.wikihow.com/Calculate-the-Volume-of-a-Square-Pyramid Look at the second method